For what values of c does the curve f(x)=2x^3+cx^2+2x have the minimum and maximum points


1
Expert's answer
2022-05-04T01:51:30-0400

"f'(x)=6x^2+2cx+2"

"f"(x)=12x+2c"

"f'(x)=0"

"6x^2+2cx+2=0"

"D=4c^2-48"

"x_1=\frac{-2c+\sqrt{4c^2-48}}{12}"

"x_2=\frac{-2c-\sqrt{4c^2-48}}{12}"

"12x(\frac{-2c+\sqrt{4c^2-48}}{12})+2c=0"

"-2c+\sqrt{4c^2-48}=-2c"

"4c^2-48=1"

"c=\sqrt{49/4}=\plusmn 3.5"


If c=3.5

"f'(x)=6x^2+7x+2=0"

"D=49-48=1"

"x=\frac{-7 \plusmn \sqrt1}{12}"

"x_1=-0.5"

"x_2=-2\/3"

"f"(-0.5)=12(-0.5)+2(3.5)=1>0"

So it is minimum

"f"(-2\/3)=12(-2\/3)+2(3.5)=-1<0"

So it is maximum


If c=-3.5

"f'(x)=6x^2-7x+2=0"

D=49-48=1

"x=\frac{7 \plusmn \sqrt1}{12}"

"x_1=0.5"

"x_2=2\/3"

"f"(-0.5)=12(0.5)+2(3.5)=13>0"

So it is minimum

f"(2/3)=12(2/3)+2(3.5)=15>0

So it is minimum